OpenStudy (anonymous):

ok whats (csc (x))^2

OpenStudy (anonymous):

1+cot^2(x)

myininaya (myininaya):

how to integrate it?

OpenStudy (anonymous):

Mine is a trig identity :P

myininaya (myininaya):

very good malevolence

OpenStudy (anonymous):

it is \[\frac{1}{\sin^2(x)}\] for one

OpenStudy (anonymous):

Haha, RECIPROCAL IDENTITY.

myininaya (myininaya):

omg satellite is is smart

myininaya (myininaya):

what do you to do with (cscx)^2

OpenStudy (anonymous):

if you want an identity involving cosecant start with \[\sin^2(x)+\cos^2(x)=1\] and then to get cosecant divide every term by sine

OpenStudy (anonymous):

you get \[\frac{sin^2(x)}{\sin^2(x)}+\frac{\cos^2(x)}{\sin^2(x)}=\frac{1}{\sin^2(x)}\]

OpenStudy (anonymous):

giving \[1+\cot^2(x)=\csc^2(x)\]

OpenStudy (anonymous):

a more or less useful trick to know

OpenStudy (anonymous):

hello my myininaya!

myininaya (myininaya):

hey

OpenStudy (anonymous):

so its 1/sin^2x?...

OpenStudy (anonymous):

that is just the reciprocal identity yes.

myininaya (myininaya):

will you tell me please what you want to do with (cscx)^2 because there is alot ways you can write that

OpenStudy (anonymous):

not sure what you are after

OpenStudy (anonymous):

almost time for coveted karnak award here.

myininaya (myininaya):

\[\int\limits_{}^{}-\csc^2xdx=cotx+C\] so what do you think the antiderivate of csc^2x is?

OpenStudy (anonymous):

oh lorda mercy. this is ok!

OpenStudy (anonymous):

look in the book and see what has derivative \[csc^2(x)\]

OpenStudy (anonymous):

i give you a hint: it is almost \[\cot(x)\]

myininaya (myininaya):

\[\frac{d}{dx}(-cotx)=\frac{d}{dx}(\frac{-cosx}{sinx})=\frac{-(-sinx)*sinx-cosx*(-cosx)}{\sin^2x}\] \[=\frac{\sin^2x+\cos^2x}{\sin^2x}=\frac{1}{\sin^2x}=\csc^2x\]

myininaya (myininaya):

so we have \[\frac{d}{dx}(-cotx)=\csc^2x\] what happens if we integrate both sides \[\int\limits_{}^{}\frac{d}{dx}(-cotx)dx=\int\limits_{}^{}\csc^2x dx\] \[-cotx+C=\int\limits_{}^{}\csc^2x dx\]

OpenStudy (anonymous):

im sorry i was occupied by my baby boy...but yes i see -cot would be -cosx/sinx

OpenStudy (anonymous):

ok so overall it would be 1/sinx^2 +3e^t^2?

myininaya (myininaya):

? no